OpenStudy (anonymous):

What is a common factor for 9, 8, 6, and 5?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[\color{red}{\Huge{\text{one !! }}}\]

myininaya (myininaya):

uno

myininaya (myininaya):

un

OpenStudy (anonymous):

Find the prime factorizations of each:\[9=3^2\]\[8=2^3\]\[6=2^1\cdot 3^1\]\[5=5^1\]There are no prime numbers that are common for all five integers, so thus the answer is 1 (the implied factor for all numbers).